500 J of work are done on a system in a process that decreases the system's thermal energy by 200 J. How much heat energy is transfered to or from the system
Answer #1
Concepts and reason

The concept used to solve this problem is the first law of thermodynamics.

We can use the first law of thermodynamics to solve it.

Fundamentals

The first law of thermodynamics states that the change in internal energy of a closed system is equal to the difference of heat applied to the system and the work done by the system.

The expression for the thermodynamic first law is,

AU=Q+W

Here, is change in internal energy, is heat, and is the work done

The internal energy is equal to the heat applied to the system minus work done by the system

AU=Q+W

Expression for the internal energy of the system is,

AU=Q+W

Substitute for and -200
for

TOOS +Õ = 1007-

Rearrange to get heat transferred from the system,

Q=-700J

Ans:

The heat transferred is COOL-
.